alyssaplaisance:

2, 11, 20, 29, 38, 47, 56,... Jane is given this sequence and is asked to write the definition. Which definition describes the sequence. A) You add 9 to every term to find the sequence. B) The first term is 2 and every term after the first is 9 more than the first term. C) The first term is 2 and every term after the first is 9 more than the preceding term. D) The first term is 2 and every term after the first is 7 more than the preceding term. E) The first term is 56 and every term after the first term is 9 less than the preceding term.

Mangaartist:

A?

Cosmicalamity:

its a because 2+9=11+9=20 and so on

iphoneking13:

its A

alyssaplaisance:

thank youll

iphoneking13:

np

alyssaplaisance:

yea